A particle decelerate uniformly from point A to B with velocity of 20m/s in time t1,if it completely comes to rest at point c with a velocity 20m/s in time t2 if it's distance between B and c is 150m/s
1:time taken between point B and C (t1). 2:Time taken between A and B (t1). 3:the total distance cover. 4:Total time taken for the whole journey
1
Expert's answer
2018-12-11T15:33:10-0500

1:

S_2=(v_f^2-v_i^2)/2a → a=(v_f^2-v_i^2)/2S=(0-〖20〗^2)/(2∙150)=-1.33 m/s^2


v_f=v_i+at_2 → t_2=(v_f-v_i)/a=(0-20)/(-1.33)≈15 s


2:

S_1=(v_f^2-v_i^2)/2a=(〖20〗^2-〖50〗^2)/(2∙(-1.33) )≈790 m


v_f=v_i+at_1 → t_1=(v_f-v_i)/a=(20-50)/(-1.33)≈23 s


3:

S_total=S_1+S_2=150+790=940 m


4:

t_total=t_1+t_2=15+23=38 s